OpenStudy (anonymous):

Which is the 25th term of the arithmetic sequence where a1 = 7 and a9 = 55 ?

OpenStudy (muhammad_nauman_umair):

AS a9=a1+8d as a1=7;a9=55 so d=(a9-a1)/8 so d=6 as a25=a1+24d a25=7+24(6) a25=151

OpenStudy (anonymous):

thanks for your help!! :)